# Geometric Reasoning in the Embedding Space   Jan Hůla1,2, David Mojžíšek2, Jiří Janeček2, David Herel1, Mikoláš Janota1

### v1

@@ Abstract
In this contribution, we demonstrate that Graph Neural Networks and Transformers can learn to reason about geometric constraints. We train them to predict spatial position of points in a discrete 2D grid from a set of constraints that uniquely describe hidden figures containing these points. Both models are able to predict the position of points and interestingly, they form the hidden figures described by the input constraints in the embedding space during the reasoning process. Our analysis shows that both models recover the grid structure during training so that the embeddings corresponding to the points within the grid organize themselves in a 2D subspace and reflect the neighborhood structure of the grid. We also show that the Graph Neural Network we design for the task performs significantly better than the Transformer and is also easier to scale.
Keywords: Machine Learning, Geometric Reasoning, Neural Networks, Embedding Space
@@ 1 Introduction
How do Neural Networks (NNs) solve problems that require spatial reasoning? Many papers already demonstrated that autoregressive Transformers, e.g. Momennejad et al. (2023); Wu et al. (2024); Yamada et al. (2023), and Graph Neural Networks (GNNs), e.g. Li et al. (2023); Teodorescu et al. (2022), are able to learn to solve such problems but it is not clear what is the mechanism they discovered.
A notable example is the work by Trinh et al. (2024) who developed a system named AlphaGeometry, which was able to solve problems that appeared on the International Mathematical Olympiad. The system contains an autoregressive Transformer that takes as input a sequence of tokens describing the problem in the language of geometric relations and is trained to predict auxiliary points useful for finding a proof for the given statement. Unfortunately, there is no explanation of the process by which the model manages to propose useful auxiliary points.
@@
In this contribution, we take a closer look at these questions in the domain of geometric Constraint Satisfaction Problems (CSPs). In order to simplify the problem as much as possible, we create a simple CSP language with several geometric constraints (relations) for which the domain is a set of points in a discrete 2D grid of certain size. Each instance of this CSP uniquely describes a hidden figure whose points are the solution of the instance. As the discrete grid is finite, we can assign a token/class [^2] When using a Transformer, each point is represented by a token; when using GNN, each point corresponds to a class. to each point and train the model to predict the points in the hidden figure with a cross-entropy loss.
Our analysis reveals interesting properties of the learned model. After visualizing a low-dimensional projection of the embeddings corresponding to individual points, we can see that they organized themselves in a 2D grid which they represent.
We also show that during inference, the embeddings of unknown variables evolve into a configuration that reflects the hidden figure described by the problem. Finally, we show that a GNN designed for the CSP performs significantly better than the autoregressive Transformer and is also able to scale to larger grids. Even though the prediction task we deal with is much simpler than the one required in systems such as AlphaGeometry, we believe that our findings provide useful insights for designing and thinking about models used for geometric reasoning.
The rest of the paper has the following structure. In Section 2, we introduce related work. Section 3 contains the description of our experimental setup. This includes the two types of architectures and the process for the generation of problems. In Section 4, we describe results of experiments for both models and provide an analysis of failure modes and scaling laws (in 4.3 and 4.4, respectively) for different sizes of the 2D grid. We test how well the described methods scale with changes in data. We discuss the results and limitations of our work in Section 5. Supplementary materials can be found in Appendix.
@@ 2 Related Work
Reasoning About Geometry
@@
Spatial Reasoning in Neural Networks
Another related area of research is focused on the ability of NNs to do spatial reasoning in a broader sense, e.g. spatial navigation, planning and interaction with physical objects.
Several papers explore spatial reasoning in the context of Large Language Models (LLMs) such as ChatGPT or LLama Touvron et al. (2023). In Momennejad et al. (2023) and Yamada et al. (2023), they point out various failure modes in a planning and navigation capability of large pool of LLMs. Wu et al. (2024) developed a prompting technique called Visualization-of-Thought (VoT), which significantly improved LLM performance in visual navigation tasks. Most similar to our work is the work by Ivanitskiy et al. (2023) who probe small Transformer models trained for maze navigation in order to understand the representation it uses to solve the task. Unlike the previously mentioned works, Janner et al. (2018) uses reinforcement learning to train a language model based on a recurrent NN (RNN) for text-based spatial reasoning.
Apart from language models, other papers also explore spatial reasoning in the context of GNNs. Li et al. (2023) developed a GNN for spatial navigation, which utilizes a memory mechanism for long-range information storage and Teodorescu et al. (2022) introduced the SpatialSim benchmark together with a GNN that is trained to recognize spatial configurations. Our work contributes into this area by partially elucidating the mechanism by which NNs are able to reason about geometric relations and by showing that GNNs which use the graph of relations/constraints and variables are much easier to train and scale.
3 Experimental Setup
@@ 3.1 Data Generation
Our goal was to create the simplest possible synthetic dataset that would enable us to study the mechanism which NNs use to reason about geometric relations. We wanted to know whether the geometry of embeddings within the network somehow reflects the geometric properties of the problem. In the task of auxiliary point prediction for which AlphaGeometry was using a language model, it is, for example, not clear whether the model is taking the geometric nature of the constraints into account or treats the language of the problem abstractly without any metric interpretation/grounding.
@@
Therefore, we designed a generator for CSPs whose solutions can be interpreted as hidden figures described by the input constraints. The task of the model is to predict positions of the points mentioned in the constraints. We simplify the problem so that the hidden figures lie on a discrete grid, and we can therefore treat the prediction of the positions as a classification task. We note that one can also treat the task as a regression and train the model to process figures with arbitrary 2D coordinates.
@@
Our CSP is using four types of constraints: M, R, S and T. The constraint $M(a,b,c)$ says that the point $b$ is a midpoint of the line segment given by $a$ and $c$ and $R(a,b,c,d)$ says that points $a$, $b$ give an axis of symmetry around which $c$ and $d$ reflect each other. The constraint $S(a,b,c,d)$ says that points $a$, $b$, $c$, $d$ form a square and $T(a,b,c,d)$ says that the vector $d-c$ is a translation of the vector $b-a$. In order to have unique solutions, we also need to fix several points to concrete positions. We can achieve that by additional constraint P for which $P(a,[1,1])$ signifies that point $a$ lies at position $[1,1]$. For each constraint, certain number of variables are required to be known so that the constraint can be uniquely resolved, here we call them determining variables. Once these variables are fixed, the rest of variables is uniquely determined. These are called dependent variables. For the constraint $M$, $2$ variables are determining and $1$ is dependent, i.e. if we know positions of $2$ variables, the last one could be resolved. For the constraint $R$, $3$ variables are determining and $1$ is dependent. For the constraint $S$, $2$ variables are determining and $2$ are dependent (i.e., we assume a fixed spatial ordering of the square vertices). For the constraint $T$, $3$ variables are determining and $1$ is dependent. The last subplot in Figure 2 shows a visualization of the following instance of the CSP:
$T(F,D,E,G)\wedge T(C,G,A,H)\ \wedge$
$S(H,I,G,J)\wedge S(B,J,L,K)\ \wedge$
$P(A,[6,15])\wedge P(B,[12,16])\wedge P(C,[12,14])\ \wedge$
$P(D,[14,12])\wedge P(E,[12,14])\wedge P(F,[12,9])\ ,$
where $A-F$ are known points and $G-L$ are unknown.
@@
We generate the problem instances in such a way that the correct assignment to all variables can be obtained by incrementally resolving constraints that can be uniquely resolved. Therefore, to sample a problem, we sample a random directed acyclic graph (DAG) with nodes corresponding to types of constraints and edges corresponding to dependencies. Then we add variables to the constraints so that each constraint within graph can be uniquely resolved given that the parent constrains are already resolved (the parent constraints therefore need to contain the determining variables). To produce a problem with a unique solution, we fix the required number of variables appearing in each of the root constraints which have no parents (i.e. arbitrary 2 variables in the S constraint, 3 variables for the T constraint, etc.) We use the Z3 solver De Moura and Bjørner (2008) to obtain the assignment for all other points. If the resulting figure is larger than the size of the grid, we reject the problem, otherwise we place it to a random position within the grid. The constraints with the fixed points are given as an input to the model, the positions of the remaining variables are given as labels.
@@ 3.2 Models
The architecture of the GNN is inspired by the model presented in Hůla et al. (2024). It is a GNN operating on a bipartite graph with $n$ variables and $m$ constraints defined by Equations (1) and (2) below which are applied recurrently for several rounds. Equation (1) updates the embeddings of variables which are, after the last update, used to classify a given variable to a point within the grid. This is achieved by applying a linear layer $L(\cdot)$ on each variable embedding. As mentioned in 3.1, some variables need to be fixed to concrete points so that the problem has a unique solution. The embeddings of these variables are initialized with the embedding layer whose weights are shared with the final classification layer $L(\cdot)$ [^3] Similarly as token embeddings share weights with the classification head in language models.. They are also not updated during message passing. The embeddings of variables that are not fixed are initialized by random unit vectors and updated after each message passing step. The update equations have the following form:
$X^{t+1}$ $=U_{X}\left(A_{X}Z^{t},X^{t},\Phi_{X}\right)\,,$ (1)
$Z^{t}_{c}$ $=U_{c}\left(A_{c}X^{t},Z^{t-1},\Phi_{c}\right)$ $\forall c\in C\,,$ (2)
@@
where $X^{t}\in\mathbb{R}^{n\times d}$, $Z^{t}\in\mathbb{R}^{m\times d}$, $A_{X}\in\mathbb{R}^{n\times m}$, $A_{c}\in\mathbb{R}^{4m\times n}$ are matrices and $C=\{M,R,S,T\}$ is the set of constraint types. Equation (1) updates the embeddings of variables and Equation (2) updates the embeddings of the constraints. The latter equation is further indexed by $c$ which reflects the fact that we are using multiple types of constraints which require distinct update functions. The update functions $U_{x}$ and $U_{c}$ are in our case realized as LSTMs (Hochreiter, 1997) parameterized by $\Phi_{X}$ and $\Phi_{c}$, respectively. Technically, each LSTM also updates a cell state which is not mentioned in the equations. These LSTMs can be also replaced by simple RNNs but we found the LSTMs easier to optimize.
@@
The update of each embedding happens independently of the other embeddings, i.e. in Equation (1), each embedding of a variable (row in the matrix $X^{t}$) is updated independently by the same LSTM which takes as input the aggregated message from the constraints containing this particular variable. The aggregated message is simply the sum of the relevant constraint embeddings and it is realized by multiplying the constraint embedding matrix $Z^{t}$ by the matrix $A_{X}$. The LSTM $U_{c}$ which updates the constraint embeddings, differs by the fact that the aggregated message is not obtained as a sum but as a concatenation of variable embeddings. The embeddings are concatenated in the order in which the variables appeared within the constraints. For example, for the constraint $S(a,b,c,d)$, we concatenate the embeddings of variables $a,b,c,d$ in that order. One can intuitively view the embeddings of variables as if they are representing the values of these variables (points) and embeddings of constraints as if they represent the information of what is needed to satisfy the constraint. For this reason, the function which updates the constraint embeddings cannot be permutation invariant because different order of the determining variables results in different values for the dependent variables. The embdding dimension $d$ is set to $96$ for the experiment in 4.2 and $256$ for the experiment in 4.4. The number of iterations of the model is set to $20$. Other hyperparameters for training the GNN can be found in Table 3.
@@
Table 1: Performance of the GNN when allowing different types of geometric constraints in problems. We report both point-wise accuracy (proportion of correctly predicted points) and complete accuracy (proportion of problems where all points were correctly predicted). Each row shows results for different combination of constraint types: SQUARE (S), TRANSLATION (T), MIDPOINT (M) and REFLECTION (R). The distribution of sampling a each type of constraint is uniform. All problems are for a grid size $20\times 20$.
Point Accuracy Complete Accuracy
S and T 98.4% 84.7%
S, T and M 89.5% 71.8%
S, T, M and R 67.7% 38%
4 Experiments

### v2

@@ Abstract
While neural networks can solve complex geometric problems, as demonstrated by systems like AlphaGeometry, we have limited understanding of how they internally represent and reason about spatial relationships. In this work, we investigate how neural networks develop internal spatial understanding by training Graph Neural Networks and Transformers to predict point positions on a discrete 2D grid from geometric constraints that describe hidden figures. We show that both models develop interpretable internal representations that mirror the geometric structure of the problems they solve. Specifically, we observe that point embeddings self-organize into 2D grid structures during training, and during inference, the models iteratively construct the hidden geometric figures within their embedding spaces. Our analysis reveals how reasoning complexity correlates with prediction accuracy, and shows that models solve constraints through an iterative refinement process, which might resemble continuous optimization. We also find that Graph Neural Networks prove more suitable than Transformers for this type of structured constraint reasoning and scale more effectively to larger problems. These findings provide initial insights into how neural networks can develop structured understanding and contribute to interpretability of neural networks.
Published Version: This paper has been published in Machine Learning and Knowledge Extraction (MAKE), 2025, 7(3): 93.
DOI: https://doi.org/10.3390/make7030093
Citation: Mojžíšek, D.; Hůla, J.; Janeček, J.; Herel, D.; Janota, M. Geometric Reasoning in the Embedding Space. Mach. Learn. Knowl. Extr. 2025, 7(3), 93.
Keywords: machine learning; geometric reasoning; neural networks; interpretability
@@ 1 Introduction
While neural networks can solve complex spatial reasoning problems, we have limited understanding of the internal mechanisms they use to represent and manipulate geometric relationships. Many papers already demonstrated that autoregressive Transformers, for example Momennejad et al. (2023); Wu et al. (2024); Yamada et al. (2023), and Graph Neural Networks (GNNs), for example Li et al. (2023); Teodorescu et al. (2022), are able to learn to solve such problems but it is not clear what mechanism they discovered.
A notable example is the work by Trinh et al. (2024) who developed a system named AlphaGeometry, which was able to solve problems that appeared on the International Mathematical Olympiad. The system contains an autoregressive Transformer that takes as input a sequence of tokens describing the problem in the language of geometric relations and is trained to predict auxiliary points useful for finding a proof for the given statement. AlphaGeometry achieves impressive performance, it operates as a black box, providing no insights into how the model internally represents geometric relationships or constructs spatial understanding.
@@
In this contribution, we take a closer look at these questions using a simplified approach. We will focus on purpose-built geometric Constraint Satisfaction Problems (CSPs). This setting allows for detailed analysis of neural geometric reasoning that is difficult to achieve in complex systems like AlphaGeometry. We create a simple CSP language with several geometric constraints (relations) for which the domain is a set of points in a discrete 2D grid of certain size. Each instance of this CSP uniquely describes a hidden figure whose points are the solution of the instance. As the discrete grid is finite, we can assign a token/class  [^2] When using a Transformer, each point is represented by a token; when using GNN, each point corresponds to a class. to each point and train the model to predict the points in the hidden figure with a cross-entropy loss.
Analysis shows that both models develop interpretable internal representations that mirror the geometric structure of the problems they solve. After visualizing a low-dimensional projection of the embeddings corresponding to individual points, we can see that they organized themselves in a 2D grid which they represent. We also show that during inference, the embeddings of unknown variables evolve into a configuration that reflects the hidden figure described by the problem. Additionally, we find that GNNs prove more suitable than Transformers for structured constraint reasoning tasks and scale more effectively to larger problems.
This controlled approach provides solid mechanistic insights into how neural networks develop spatial understanding and contributes to broader understanding of structured reasoning in neural systems.
The rest of the paper has the following structure. In Section 2, we review related work. Section 3 describes our experimental setup, including the two types of architectures and the process for generating problems. In Section 4, we present results of experiments for both models, analyze embedding structure emergence, examine the iterative solution process, and provide analysis of failure modes for different problem complexities. We discuss the results and limitations of our work in Section 5. Supplementary materials can be found in Appendix.
@@ 2 Related Work
2.1 Reasoning About Geometry
@@
Other neural approaches include visual reasoning methods that extract geometric primitives from diagrams using deep learning for construction problems Wong et al. (2022). Our work differs by focusing on abstract constraint satisfaction rather than visual diagram processing, enabling detailed analysis of internal spatial representations.
@@
2.2 Spatial Reasoning in Neural Networks
Another related area of research is focused on the ability of NNs to do spatial reasoning in a broader sense, for example spatial navigation, planning and interaction with physical objects.
Several papers explore spatial reasoning in the context of Large Language Models (LLMs) such as ChatGPT or LLama Touvron et al. (2023). In Momennejad et al. (2023) and Yamada et al. (2023), they point out various failure modes in a planning and navigation capability of large pool of LLMs. In Wu et al. (2024) they developed a prompting technique called Visualization-of-Thought (VoT), which significantly improved LLM performance in visual navigation tasks. Most similar to our work is the work by Ivanitskiy et al. (2023) who probe small Transformer models trained for maze navigation in order to understand the representation it uses to solve the task. Unlike the previously mentioned works, Janner et al. (2018) uses reinforcement learning to train a language model based on a recurrent NN (RNN) for text-based spatial reasoning.
Apart from language models, other papers also explore spatial reasoning in the context of GNNs. In Li et al. (2023) they developed a GNN for spatial navigation, which utilizes a memory mechanism for long-range information storage and Teodorescu et al. (2022) introduced the SpatialSim benchmark together with a GNN that is trained to recognize spatial configurations. Our work contributes into this area by partially elucidating the mechanism by which NNs are able to reason about geometric relations and by showing that GNNs which use the graph of relations/constraints and variables are much easier to train and scale.
3 Problems and Methods
@@ 3.1 Geometric Problem Generation
Our goal was to create a simple synthetic dataset that would enable us to study the mechanism which NNs use to reason about geometric relations. Specifically, we wanted to investigate whether neural networks develop internal representations that reflect the geometric structure of the problems they solve. In the task of auxiliary point prediction for which AlphaGeometry was using a language model, it is, for example, not clear whether the model is taking the geometric nature of the constraints into account or treats the language of the problem abstractly without any metric interpretation/grounding.
@@
Therefore, we designed a generator for CSPs whose solutions can be interpreted as geometric figures described by the input constraints. The task of the model is to predict positions of the points contained in the constraints. We simplify the problem so that the geometric figures lie on a discrete grid, and we can therefore treat the prediction of the positions as a classification task. We note that one can also treat the task as a regression and train the model to process figures with arbitrary 2D coordinates.
@@
Our CSP is using four types of constraints: $M$, $R$, $S$ and $T$. The constraint $M(A,B,C)$ says that the point $B$ is a midpoint of the line segment given by $A$ and $C$ and $R(A,B,C,D)$ says that points $A$, $B$ give an axis of symmetry around which $C$ and $D$ reflect each other. The constraint $S(A,B,C,D)$ says that points $A$, $B$, $C$, $D$ form a square and $T(A,B,C,D)$ says that the vector DCD-C is a translation of the vector BAB-A. In order to have unique solutions, we also need to fix several points to concrete positions. We can achieve that by additional constraint $P$ for which $P(a,[1,1])$ signifies that point $A$ lies at position $[1,1]$. For each constraint, certain number of variables are required to be known so that the constraint can be uniquely resolved, here we call them determining variables. Once these variables are fixed, the rest of variables is uniquely determined. These are called dependent variables. For the constraint $M$, 22 variables are determining and $1$ is dependent, i.e. if we know positions of $2$ variables, the last one could be resolved. For the constraint $R$, 33 variables are determining and $1$ is dependent. For the constraint $S$, 22 variables are determining and 22 are dependent (i.e., we assume a fixed spatial ordering of the square vertices). For the constraint $T$, 33 variables are determining and $1$ is dependent. The last subplot in Figure 7 shows a visualization of the following instance of the CSP:
$T(F,D,E,G)$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $T(C,G,A,H)$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $S(H,I,G,J)$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $S(B,J,L,K)$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $P(A,[6,15])$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$
$P(B,[12,16])$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $P(C,[12,14])$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $P(D,[14,12])$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $P(E,[12,14])$ $\mathrel{\hskip-5.0pt\wedge\hskip-5.0pt}$ $P(F,[12,9])$ $\ \ \ \ \ ,$
where AFA-F are known points and GLG-L are unknown.
Our generator creates problems that require deductive reasoning by ensuring constraints form a dependency structure where some must be resolved before others. Attempting to solve all constraints simultaneously is not feasible because dependent variables would be incorrectly determined. For example, a network might successfully create a square from four unknown points, but if some of those points are actually determined by previous translation or reflection constraints, the resulting square would be geometrically valid but positioned incorrectly within the overall figure. This dependency structure forces models to discover the proper reasoning sequence rather than solving constraints in isolation.
@@
To create a single problem, we sample a random directed acyclic graph (DAG) with nodes corresponding to types of constraints and edges corresponding to dependencies. Then we add variables to the constraints so that each constraint within graph can be uniquely resolved given that the parent constraints are already resolved (the parent constraints therefore need to contain the determining variables). To produce a problem with a unique solution, we fix the required number of variables appearing in each of the root constraints which have no parents (i.e. arbitrary two variables in the S constraint, three variables for the T constraint, etc.) We use the Z3 solver De Moura and Bjørner (2008) to obtain the assignment for all other points. If the resulting figure is larger than the size of the grid, we reject the problem, otherwise we place it to a random position within the grid. The constraints with the fixed points are given as an input to the model, the positions of the remaining variables are given as labels.
By varying the number and types of constraints, the generator produces problems with different complexity levels and reasoning depths. This allows us to investigate whether models can solve problems harder than those seen during training by leveraging test-time scaling approaches such as increased iterations or resampling.
@@ 3.2 Models
The architecture of the GNN is inspired by the model presented in Hůla et al. (2024) and originated in Selsam et al. (2018). Similarly as in works about Boolean satisfiability, it applies the same update rules repeatedly. It is a GNN operating on a bipartite graph with $n$ variables and $m$ constraints defined by Equations (2) and (1) below which are applied recurrently for several rounds. Equation (2) updates the embeddings of variables which are, after the last update, used to classify a given variable to a point within the grid. This is achieved by applying a linear layer $L(\cdot)$ on each variable embedding. As mentioned in Section 3.1, some variables need to be fixed to concrete points so that the problem has a unique solution.
The embeddings of known variables are initialized using the embedding layer that shares weights with the final classification layer $L(\cdot)$, i.e., both use the same, trainable, weight matrix $\mathbf{W}\in\mathbb{R}^{N\times d}$, where $N$ is the number of grid positions and $d$ is the embedding dimension. This design ensures consistent representation of grid positions throughout the network. [^3] Similarly as token embeddings share weights with the classification head in language models. They are not updated during message passing. The embeddings of unknown variables are initialized randomly, i.e., they are sampled from a unit hypercube $0,1]^{d}$ following the uniform probability distribution of each coordinate. It enables resampling during inference by using different random initializations.
The embeddings are updated after each message passing iteration. Let $n$ be the number of variables, $m$ be the number of constraints. The update equations have the following form:
$Z^{t+1}_{c}$ $=U_{c}\left(\text{cat}\left(A_{c}X^{t}\right),Z^{t}_{c},\Phi_{c}\right)$ $\forall c\in C\,$, (1)
$X^{t+1}$ $=U_{X}\left(A_{X}Z^{t+1},X^{t},\Phi_{X}\right)\,$, (2)
[…]
@@
Equation (2) updates the embeddings of variables and Equation (1) updates the embeddings of the constraints. The latter equation is further indexed by $c$ which reflects the fact that we are using multiple types of constraints which require distinct update functions. The update functions $U_{x}$ and $U_{c}$ ($\forall c\in C$) are in our case realized as LSTMs (Hochreiter, 1997) parameterized by $\Phi_{X}$ and $\Phi_{c}$, respectively. Technically, each LSTM also updates a cell state which is not mentioned in the equations. These LSTMs could be replaced by simple RNNs but we found the LSTMs easier to optimize and RNN does not work well with our data, see A.4. The simplified schema of the GNN processing an example can be found in Figure 1.
Figure 1: Solution process of the GNN model. The geometric problem is encoded as a bipartite graph with variable nodes (points) and constraint nodes. Red points represent fixed/known variables that maintain their initial embeddings from the embedding layer throughout the process. Green points represent unknown variables and blue nodes represent constraint embeddings, both initialized randomly. The embedding layer is represented as a matrix with dimensions NN (corresponding to grid points i.e. classes) by $d$ (embedding dimension, $d=128$). During inference, unknown point embeddings and constraint embeddings are iteratively updated using respective LSTMs: $U_{x}$ updates variable embeddings (Equation 2) and $U_{c}$ updates constraint embeddings (Equation 1) for a specified number of iterations. After the final update, unknown points are classified using a linear classification layer without bias that shares weights with the embedding layer. The classifier produces logits that are converted to probabilities via softmax to predict the final point positions.
@@
The update of each embedding happens independently of the other embeddings, i.e. in Equation (2), each embedding of a variable (row in the matrix $X^{t}$) is updated independently by the same LSTM which takes as input the aggregated message from the constraints containing this particular variable. The aggregated message is simply the sum of the relevant constraint embeddings and it is realized by multiplying the constraint embedding matrix $Z^{t}$ by the matrix $A_{X}$. The LSTM $U_{c}$ which updates the constraint embeddings, differs by the fact that the aggregated message is not obtained as a sum but as a concatenation of variable embeddings. The embeddings are concatenated in the order in which the variables appeared within the constraints. For example, for the constraint $S(A,B,C,D)$, we concatenate the embeddings of variables $A,B,C,D$ in that order. One can intuitively view the embeddings of variables as if they are representing the values of these variables (points) and embeddings of constraints as if they represent the information of what is needed to satisfy the constraint. For this reason, the function which updates the constraint embeddings cannot be permutation invariant because different order of the determining variables results in different values for the dependent variables.
The embedding dimension $d$ and number of iterations are configurable parameters that we adjust based on problem complexity. Importantly, since the same update rules are applied recurrently, the number of iterations can be modified at test time without changing the model parameters. This architectural property, leads to the option of allocating more computation to solve harder problems during inference. Other hyperparameters for training the GNN can be found in Table 2.
For clarity, one illustrative example of how a GNN processes and solves a specific problem will be presented later in 3.3.
@@
3.3 GNN: Illustrative Example
Figure 2: Graph encoding of the illustrative example problem showing the bipartite structure with point nodes and constraint nodes, along with the point-to-index mapping for the 20×20 grid.
To illustrate how the GNN works, we present a concrete example showing how it processes geometric constraints on a $N=20\times 20$ grid with embedding dimension $d=128$ and iteration count $t=15$. Consider the following constraint satisfaction problem:
$T(A,B,C,D)\land S(B,D,E,F)\land S(B,F,G,H)\land$
$P(A,[5,2])\land P(B,[2,2])\land P(C,[4,3])$
where $A,B,C$ are known points fixed by position constraints $P$, and $D,E,F,G,H$ are unknown points to be predicted.
Problem Encoding and Dependencies
Each grid position $(x,y)$ maps to a unique index using $\text{index}=x+y\times 20$. Therefore, the known points correspond to indices: $A\rightarrow 45$, $B\rightarrow 42$, $C\rightarrow 64$. The problem forms a bipartite graph with 8 point nodes and 3 constraint nodes, where the dependency structure requires sequential resolution: constraint $T_{1}$ determines point $D$, which enables constraint $S_{1}$ to determine points $E$ and $F$, which finally allows constraint $S_{2}$ to determine points $G$ and $H$.
The point constraints $P$ are not realized as separate constraint nodes in the bipartite graph, but rather as fixed embeddings taken directly from the embedding layer. Figure 2 illustrates this encoding and the target solution.
Embedding Initialization
The model uses a shared embedding matrix $\mathbf{W}\in\mathbb{R}^{400\times 128}$ for both the embedding layer and final classifier. Known points are initialized with fixed embeddings $\mathbf{W}_{45}$, $\mathbf{W}_{42}$, $\mathbf{W}_{64}\in\mathbb{R}^{128}$ corresponding to points $A$, $B$, $C$ respectively, where $\mathbf{W}_{i}$ denotes the $i$-th row of matrix $\mathbf{W}$. These embeddings remain unchanged throughout inference:
$\mathbf{x}_{A}^{(t)}=\mathbf{W}_{45},\qquad\mathbf{x}_{B}^{(t)}=\mathbf{W}_{42},\qquad\mathbf{x}_{C}^{(t)}=\mathbf{W}_{64}\quad\text{for all }t.$
Unknown points are initialized with random vectors $\mathbf{r}_{D},\mathbf{r}_{E},\mathbf{r}_{F},\mathbf{r}_{G},\mathbf{r}_{H}\sim\mathcal{U}(0,1)^{128}$, sampled uniformly from the unit hypercube, and constraint embeddings are initialized as random vectors $\mathbf{c}_{1},\mathbf{c}_{2},\mathbf{c}_{3}\sim\mathcal{U}(0,1)^{128}$. These random vectors are used as the initial hidden states for the LSTM-based message passing. Specifically, for each unknown point $i$, we set the initial embedding as the hidden state $\mathbf{x}_{i}^{(0)}:=\mathbf{r}_{i}$, and similarly for each constraint $c$, we set $\mathbf{z}_{c}^{(0)}:=\mathbf{c}_{c}$. The LSTM cell states $\mathbf{h}_{i}^{(0)}$ and $\mathbf{h}_{c}^{(0)}$ are initialized to zero. During inference, the hidden states $\mathbf{x}_{i}^{(t)}$ and $\mathbf{z}_{c}^{(t)}$ evolve over time according to the update equations.
Message Passing Dynamics
The update equations are repeated for 15 iterations. Our LSTM-based architecture maintains both hidden states and cell states, where the hidden states serve as the embeddings that flow through the message passing network, while cell states maintain internal memory. The LSTM also produces a cell output at each iteration, but this output is discarded and only the hidden and cell states are retained.
For constraint updates (Equation 1), each constraint type uses a specialized LSTM update function. The constraint embedding $\mathbf{z}_{c}^{(t)}$ (LSTM hidden state) and cell state $\mathbf{h}_{c}^{(t)}$ are updated as:
$\mathbf{z}_{c}^{(t+1)},\mathbf{h}_{c}^{(t+1)}=U_{c}^{(\text{type})}(\mathbf{m}_{c}^{(t)},\mathbf{z}_{c}^{(t)},\mathbf{h}_{c}^{(t)})$
where $\mathbf{m}_{c}^{(t)}$ is the concatenated message from variables participating in constraint $c$, and each constraint type has its own LSTM parameters: $U_{c}^{(T)}$ for translation, $U_{c}^{(S)}$ for square, $U_{c}^{(R)}$ for reflection, and $U_{c}^{(M)}$ for midpoint constraints. The LSTM returns both an output (which is discarded) and a tuple containing the new hidden state $\mathbf{z}_{c}^{(t+1)}$ and cell state $\mathbf{h}_{c}^{(t+1)}$.
For the above example, the input message $\mathbf{m}_{T_{1}}^{(t)}$ for constraint $T_{1}$ is the concatenated vector $\mathbf{W}_{45};\mathbf{W}_{42};\mathbf{W}_{64};\mathbf{x}_{D}^{(t)}]\in\mathbb{R}^{512}$, where $\mathbf{W}_{45}$, $\mathbf{W}_{42}$, and $\mathbf{W}_{64}$ are the fixed embeddings of known points $A$, $B$, and $C$ respectively, and $\mathbf{x}_{D}^{(t)}$ is the current hidden state of point $D$. Similarly, the messages $\mathbf{m}_{S_{1}}^{(t)}$ and $\mathbf{m}_{S_{2}}^{(t)}$ for constraints $S_{1}$ and $S_{2}$ are given by $\mathbf{W}_{42};\mathbf{x}_{D}^{(t)};\mathbf{x}_{E}^{(t)};\mathbf{x}_{F}^{(t)}]$ and $\mathbf{W}_{42};\mathbf{x}_{F}^{(t)};\mathbf{x}_{G}^{(t)};\mathbf{x}_{H}^{(t)}]$ respectively. Each of these message vectors is passed to the corresponding constraint-specific LSTM update function $U_{c}^{(\text{type})}$.
For variable updates (Equation 2), unknown point embeddings are updated using messages aggregated from all connected constraints. Each unknown variable $i$ aggregates the sum of constraint embeddings from all connected constraints and passes this sum to the shared LSTM update function $U_{X}$:
$\mathbf{x}_{i}^{(t+1)},\mathbf{h}_{i}^{(t+1)}=U_{X}\left(\sum_{c\in\mathcal{N}(i)}\mathbf{z}_{c}^{(t+1)},\mathbf{x}_{i}^{(t)},\mathbf{h}_{i}^{(t)}\right)$
Here, $\mathcal{N}(i)$ denotes the set of constraint nodes connected to variable $i$, $\mathbf{x}_{i}^{(t)}$ is the hidden state (embedding) of point $i$ at time step $t$, and $\mathbf{h}_{i}^{(t)}$ is its corresponding LSTM cell state.
In the illustrative example, point $D$ receives messages from constraints $T_{1}$ and $S_{1}$, resulting in an update input of $\mathbf{z}_{T_{1}}^{(t+1)}+\mathbf{z}_{S_{1}}^{(t+1)}$. Point $F$ receives messages from $S_{1}$ and $S_{2}$. Points $E$, $G$, and $H$ each receive a single constraint message: from $S_{1}$ for $E$, and from $S_{2}$ for $G$ and $H$.
These aggregated messages are passed into the $U_{X}$ LSTM, which outputs the updated hidden state $\mathbf{x}_{i}^{(t+1)}$ that becomes the new embedding of variable $i$.
Solution Recovery
After $t$ iterations, the model predicts the final grid position of each unknown variable using a standard classification head. The final embedding $\mathbf{x}_{i}^{(T)}$ is projected using the shared embedding matrix $\mathbf{W}\in\mathbb{R}^{400\times 128}$ (no bias term) to produce logits:
$\boldsymbol{\ell}_{i}=\mathbf{W}\mathbf{x}_{i}^{(T)}\in\mathbb{R}^{400}$
The predicted index is obtained as $\hat{k}_{i}=\arg\max(\boldsymbol{\ell}_{i})$, which is then mapped back to coordinates via $\hat{x}=\hat{k}_{i}\bmod 20$, $\hat{y}=\lfloor\hat{k}_{i}/20\rfloor$.
During training, we apply a cross-entropy loss over the logits $\boldsymbol{\ell}_{i}$, comparing predicted and ground-truth indices for all unknown variables. The embedding matrix $\mathbf{W}$ is shared with the embedding layer used for known variables, ensuring consistency between input encoding and prediction.
In our example, the correct predictions for points DDHH are $[1,3]$, $[0,2]$, $[1,1]$, $[2,0]$, and $[3,1]$, satisfying all constraints.
4 Experiments and Results
We evaluate both architectures on geometric constraint problems. We examine prediction accuracy, dataset complexity effects, embedding structure emergence, initialization strategies, solution dynamics, and failure modes.